# DiffPO: Diffusion-styled Preference Optimization for Efficient Inference-Time Alignment of Large Language Models

### v2

@@ 5 Related Works
Existing training-based or inference-time alignment approaches typically rely on token-level rewards, while human preferences are generally provided and defined at the sentence level [40, 2, 69]. To address this discrepancy, some works [41, 64, 69] leverage token-wise or step-wise information to improve alignment performance. In contrast, this paper proposes modeling alignment as a sentence-level denoising process. We introduce a model-agnostic, inference-time alignment method, and our empirical results demonstrate its superiority in both performance and efficiency.
5.2 Parallel Decoding and Diffusion Process.
@@
Parallel decoding has been increasingly utilized and developed in recent research to accelerate the inference processes of large language models (LLMs). One line of research, including works by Leviathan et al. [37], Chen et al. [11], focuses on speculative decoding. These techniques enhance LLM decoding speed by employing a smaller draft model to predict the outputs, which are then verified in parallel by a larger target model. Another research trajectory explores parallel decoding strategies that do not rely on a draft model. Methods such as conditioning on “look-ahead” tokens or employing Jacobi iterations have been investigated by Santilli et al. [51], Fu et al. [23]. These approaches allow the target model to produce several tokens simultaneously, aiming for rapid convergence to a fixed point on a Jacobi trajectory. CLLMs [57] develop a novel approach, refining the target LLM to consistently predict the fixed point from any given state.
@@
Diffusion models have demonstrated significant diversity and controllability in visual generation [29, 56, 19, 12, 8, 7, 33, 6, 17, 18, 48, 68]. Recently, these models have been extended to text generation, as evidenced by the works of [38, 24, 42]. In essence, diffusion models execute a multi-step denoising process that progressively transforms random noise into a coherent data sample. In the context of text, diffusion models can be considered an evolution of traditional iterative Non-Autoregressive models, as described by Gong et al. (2022). These models have demonstrated the ability to match or surpass Autoregressive (AR) models in terms of text perplexity [28, 26], diversity [25, 70], and various sequence-to-sequence tasks [67, 66].
@@
In this paper, we are motivated by the goal of aligning Large Language Models (LLMs) with human values or intentions, as outlined in [65]. We define preferences at the sentence-level, focusing on the style or format of complete answers generated by the LLMs. If we consider each iteration of parallel decoding as a transition between states, this bears a formal resemblance to discrete diffusion models. In DiffPO, we leverage parallel decoding to implement sentence-level denoising, thereby enhancing the modeling of the alignment process.
The development of DiffPO is also inspired by Consistency Models [57] and CLLMs [35]. Consistency models address the limitation of the slow iterative sampling process by mapping any point along the probability flow ODE of the diffusion process back to the original point in a single step. CLLMs propose accelerating LLM inference by mapping the intermediate process of LLM parallel decoding to the final process. Similar to these works, we optimize DiffPO with consistency loss, thus enabling model-agnostic alignment.
6 Experiment
6.1 Experimental Setups
@@
As for the training set, we collect 6 generations from 6 base models (i.e., Llama-3-8B-Instruct, Llama-3-8B-SFT, Mistral-7B-SFT, Mistral-7B-Instruct, Gemma-2-2B-Instruct, Gemma-2-9B-Instruct). We then employ ArmoRM [61] to score these responses. The response with the highest score is selected as $\mathbf{y}^{(T)}$. The remaining five responses are ranked according to their scores to serve as $\mathbf{y}^{(0:T-1)}$. In the training process, at each iteration, we randomly sample $\mathbf{y}^{t}$ from $\mathbf{y}^{(0:T-1)}$ for optimization. We train DiffPO models using the following hyperparameters: a learning rate of 1e-9, a batch size of 1 and gradient accumulation steps of 4, a max sequence length of 1024, and a cosine learning rate schedule with 3% warmup steps for 1 epoch. All the models are trained with an Adam optimizer. All the training experiments in this paper were conducted on 8×A100 GPUs.
@@
For the MT-bench, we use GPT-4 as the judge model, following the default settings. The scores are based on a single-answer rating scale from 1 to 10. For AlpacaEval, we use GPT-4 Turbo as the judge model, which performs pairwise comparison of responses generated by GPT-4, each with the same maximum length. For HH-RLHF, we use ArmoRM for single-answer rating and report the overall score, along with the “helpful” and “harmless” scores, which are provided in dimensions 9 and 10, respectively.
@@ 6.2 Experimental Results
7 Analysis
Loss MT AE2 HH
GPT-4 LC (%) Avg.
Vanilla Decoding
Llama-3-Instruct 6.78 36.83 0.7985
$L_{\mathrm{AR}}$ 6.90 36.35 0.7971
Llama-3-it w. DiffPO-8B
$L_{\mathrm{AR}}$ only 6.75 35.84 0.7968
$10\times L_{\mathrm{Con}}+L_{\mathrm{AR}}$ 6.85 35.96 0.7997
$100\times L_{\mathrm{Con}}+L_{\mathrm{AR}}$ 6.86 35.92 0.7998
$1,000\times L_{\mathrm{Con}}+L_{\mathrm{AR}}$ 7.02 36.44 0.7998
7.1 Loss and Hyperparameter Ablation.
We supplement the ablation results for DiffPO-8B, presented in Table 7. We report on two decoding strategies: vanilla decoding of the single base model and the optimized DiffPO-8B, and DiffPO decoding, which applies the optimized DiffPO-8B on the output of the base model. The findings indicate that applying DiffPO to the base model achieves performance superior to that of single models alone, thus demonstrating the effectiveness of the DiffPO strategy. Furthermore, we report the results of an ablation study on the hyperparameter $w$ in Eq. 6. When using DiffPO decoding, employing $L_{\mathrm{Con}}$ with larger values of $w$ lead to a more pronounced improvement in performance.
7.2 Illustration of the Speed-up of DiffPO
As shown in Figure 4, AR decoding (e.g., Aligner [31]) typically generates only one aligned token per iteration. In contrast, DiffPO enables the skipping of satisfied tokens, thereby avoiding the time latency associated with token-level generation. As a result, DiffPOcan predict the modified subsequence in 3 iterations, achieving the same result as 11 iterations of AR decoding.

### v3

@@ Appendix A Related Works
Existing training-based or inference-time alignment approaches typically rely on token-level rewards, while human preferences are generally provided and defined at the sentence level Li et al. (2023b); Ahmadian et al. (2024); Zeng et al. (2024). To address this discrepancy, some works Lightman et al. (2023); Yang et al. (2024b); Zeng et al. (2024) leverage token-wise or step-wise information to improve alignment performance. In contrast, this paper proposes modeling alignment as a sentence-level denoising process. We introduce a model-agnostic, inference-time alignment method, and our empirical results demonstrate its superiority in both performance and efficiency.
A.2 Parallel Decoding and Diffusion Process.
@@
Parallel decoding has been increasingly utilized and developed in recent research to accelerate the inference processes of large language models (LLMs). One line of research, including works by Leviathan et al. (2023); Chen et al. (2023), focuses on speculative decoding. These techniques enhance LLM decoding speed by employing a smaller draft model to predict the outputs, which are then verified in parallel by a larger target model. Another research trajectory explores parallel decoding strategies that do not rely on a draft model. Methods such as conditioning on “look-ahead” tokens or employing Jacobi iterations have been investigated by Santilli et al. (2023); Fu et al. (2024). These approaches allow the target model to produce several tokens simultaneously, aiming for rapid convergence to a fixed point on a Jacobi trajectory. CLLMs Song et al. (2023) develop a novel approach, refining the target LLM to consistently predict the fixed point from any given state.
@@
Diffusion models have demonstrated significant diversity and controllability in image generation Ho et al. (2020); Song et al. (2020); Dhariwal and Nichol (2021). Recently, these models have been extended to text generation, as evidenced by the works of Li et al. (2022); Gong et al. (2022); Lovelace et al. (2024). In essence, diffusion models execute a multi-step denoising process that progressively transforms random noise into a coherent data sample. In the context of text, diffusion models can be considered an evolution of traditional iterative Non-Autoregressive models, as described by Gong et al. (2022). These models have demonstrated the ability to match or surpass Autoregressive (AR) models in terms of text perplexity Han et al. (2022); Gulrajani and Hashimoto (2024), diversity Gong et al. (2023); Zhang et al. (2024), and various sequence-to-sequence tasks Ye et al. (2024b, a).
@@
In this paper, we are motivated by the goal of aligning Large Language Models (LLMs) with human values or intentions, as outlined in Yao et al. (2023). We define preferences at the sentence-level, focusing on the style or format of complete answers generated by the LLMs. If we consider each iteration of parallel decoding as a transition between states, this bears a formal resemblance to discrete diffusion models. In DiffPO, we leverage parallel decoding to implement sentence-level denoising, thereby enhancing the modeling of the alignment process.
The development of DiffPO is also inspired by Consistency Models Song et al. (2023) and CLLMs Kou et al. (2024). Consistency models address the limitation of the slow iterative sampling process by mapping any point along the probability flow ODE of the diffusion process back to the original point in a single step. CLLMs propose accelerating LLM inference by mapping the intermediate process of LLM parallel decoding to the final process. Similar to these works, we optimize DiffPO with consistency loss, thus enabling model-agnostic alignment.
Appendix B Experiment
B.1 Experimental Setups
@@
As for the training set, we collect 6 generations from 6 base models (i.e., Llama-3-8B-Instruct, Llama-3-8B-SFT, Mistral-7B-SFT, Mistral-7B-Instruct, Gemma-2-2B-Instruct, Gemma-2-9B-Instruct). We then employ ArmoRM Wang et al. (2024) to score these responses. The response with the highest score is selected as $\mathbf{y}^{(T)}$. The remaining five responses are ranked according to their scores to serve as $\mathbf{y}^{(0:T-1)}$. In the training process, at each iteration, we randomly sample $\mathbf{y}^{t}$ from $\mathbf{y}^{(0:T-1)}$ for optimization. We train DiffPO models using the following hyperparameters: a learning rate of 1e-9, a batch size of 1 and gradient accumulation steps of 4, a max sequence length of 1024, and a cosine learning rate schedule with 3% warmup steps for 1 epoch. All the models are trained with an Adam optimizer. All the training experiments in this paper were conducted on 8×A100 GPUs.
@@
For the MT-bench, we use GPT-4 as the judge model, following the default settings. The scores are based on a single-answer rating scale from 1 to 10. For AlpacaEval, we use GPT-4 Turbo as the judge model, which performs pairwise comparison of responses generated by GPT-4, each with the same maximum length. For HH-RLHF, we use ArmoRM for single-answer rating and report the overall score, along with the “helpful" and “harmless" scores, which are provided in dimensions 9 and 10, respectively.
@@ B.2 Experimental Results
Appendix C Analysis
C.1 Illustration of the Speed-up of DiffPO
As shown in Figure 4, AR decoding (e.g., Aligner Ji et al. (2024)) typically generates only one aligned token per iteration. In contrast, DiffPO enables the skipping of satisfied tokens, thereby avoiding the time latency associated with token-level generation. As a result, DiffPOcan predict the modified subsequence in 3 iterations, achieving the same result as 11 iterations of AR decoding.
Appendix D More Related Works
LLM Pluralism and Fairness
LLM alignment ensures AI systems follow human intentions and values Stiennon et al. (2020); Bai et al. (2022); Ouyang et al. (2022); Achiam et al. (2023). However, within a single task, users’ goals and values often differ. As AI systems are increasingly used by diverse groups, they must address a broader range of needs. In short, we need AI systems that are pluralistic and fair, being capable of reflecting diverse human values Chen et al. (2024c, b); Fan et al. (2024b); Luo et al. (2024); Fan et al. (2024a); Chen et al. (2024a).